Find the equations of the lines perpendicular to the line x – y + 3 = 0 and passing at a distance numerically twice as far from the origin.


1
Expert's answer
2021-11-02T15:03:19-0400

Given line is "x- y+3=0"

The slope of this line is "m_{1}=1"

So, the slope of the required line will be "m_{2}=-1"

Since the lines are perpendicular, the required equation is in the form of "x+y+k=0".

The perpendicular distance from the origin is 2 units.

So,

"2=\frac{|(0)+(0)+k|}{\sqrt{1+1}}
\\2=\frac{|k|}{\sqrt{2}} 
\\ |k|= 2 \sqrt{2}
\\k=\pm 2 \sqrt{2}"

Thus, the required equation of line is "x+y\pm 2 \sqrt{2}=0."